OpenStudy (anonymous):

Urgent PLEASEEEE Help! Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (−6, 3). a.y=3/2x-1 b.y=3/2x+3 c.y=-2/3x-1 d.y=-2/3x+3

OpenStudy (anonymous):

What is the slope of your line?

OpenStudy (anonymous):

y = 3/2x + 4 i think . I'm not really good with this

OpenStudy (anonymous):

your equation is in the form y=mx+b, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this website is hard to use

OpenStudy (anonymous):

ok, so... m = slope, correct?

OpenStudy (anonymous):

yes I believe so

OpenStudy (anonymous):

what is your slope?

OpenStudy (anonymous):

3/2x ?

OpenStudy (anonymous):

its 3/2... yes. soo to find the slope that is perpendicular, you should find the negative reciprical

OpenStudy (anonymous):

so -2/3x?

OpenStudy (anonymous):

yes. you need three things to find the slope of a line.... point, which you have (-6,3) slope, which you just found... m=-2/3 and a form, y-y1=m(x-x1) so... y-3=-2/3(x+6) do you see how I got that?

OpenStudy (anonymous):

three things to find the equation of a line(sorry)

OpenStudy (anonymous):

yeah I understand how you got it :)

OpenStudy (anonymous):

then you just solve for y... y-3=-2/3x -4 y=-2/3x-1

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

Thank you so much! That really helped me understand how to solve problems like that. Thank you again!

OpenStudy (anonymous):

Just remember... you need three things to find the equation of a line: point (x,y) slope of m and a form (y-the given y)=m(x-the given x) and then you just solve for y

OpenStudy (anonymous):

thanks I will remember that next time!